Sunday, 25 January 2009

Maintenance: 00:00 to 06:00 25th Jan 08

Following the air-con problem yesterday it turned out that our elderly server ITO1, which has been performing front line service since we started the company, was not well. We were aware it was pretty heat-sensitive and had already purchased and two installed new fast heat-tolerant replacements with a switch-over scheduled in the next few days. We decided to bring this forward and completed it last night (UK time). ITO1 will now go into museum service and two new servers have taken up the strain.

It is likely, however that we missed an OSM update last night for the second day running last night and are therefore still running with Thursday's OSM dataset.

Saturday, 24 January 2009

Outage 24th Jan 2009 04:00 UTC to 15:00 UTC

We had an outage today from 04:00 until about 14:00 on the 24th Jan 2009. This was caused by a air conditioning failure at our hosting company which caused problems with one of our servers which was position in a hot-spot within the server room.

We apologise for any inconvenience and assure you we take such outages of our services very seriously. This is our longest outage for well over a year and we are making a number of urgent changes to reduce the chances of a repeat performance.

In future do check for details of any outage if you cannot access our site. If you spot a problem we haven't reported then please email us at

We believe our services are now back to 100% again.